### v1

# Truthful Text Sanitization Guided by Inference Attacks

Abstract
The purpose of text sanitization is to rewrite those text spans in a document that may directly or indirectly identify an individual, to ensure they no longer disclose personal information. Text sanitization must strike a balance between preventing the leakage of personal information (privacy protection) while also retaining as much of the document’s original content as possible (utility preservation). We present an automated text sanitization strategy based on generalizations, which are more abstract (but still informative) terms that subsume the semantic content of the original text spans. The approach relies on instruction-tuned large language models (LLMs) and is divided into two stages. The LLM is first applied to obtain truth-preserving replacement candidates and rank them according to their abstraction level. Those candidates are then evaluated for their ability to protect privacy by conducting inference attacks with the LLM. Finally, the system selects the most informative replacement shown to be resistant to those attacks. As a consequence of this two-stage process, the chosen replacements effectively balance utility and privacy. We also present novel metrics to automatically evaluate these two aspects without the need to manually annotate data. Empirical results on the Text Anonymization Benchmark show that the proposed approach leads to enhanced utility, with only a marginal increase in the risk of re-identifying protected individuals compared to fully suppressing the original information. Furthermore, the selected replacements are shown to be more truth-preserving and abstractive than previous methods.
keywords:
data privacy, text sanitization , data utility , truth-preserving replacements , large language models
journal: Information Sciences
[…]
[inst1]organization=Norwegian Computing Center,addressline=Postboks 114 Blindern, city=Oslo, postcode=NO-0314, country=Norway
[…]
[inst2]organization=Department of Computer Engineering and Mathematics. CYBERCAT.,addressline=Universitat Rovira i Virgili, city=Tarragona, postcode=43007, country=Spain
1 Introduction
Text documents play a key role in many sectors, such as healthcare, legal services or public administrations. However, those documents often contain personal data that, if released without adequate protection, can pose significant privacy risks to the individuals they concern. Regulations such as the General Data Protection Regulation (GDPR) [15] impose strict constraints how personal information should be protected. In particular, such data cannot be shared with third parties or employed for secondary purposes without a proper legal ground, such as the explicit consent of the individuals the data refers to. However, if the data can be fully anonymized, those restrictions no longer apply, as the resulting data is by definition no longer considered personal. Many approaches have been proposed to perform this anonymization [24], although their compliance relative to legal frameworks such as GDPR remains unclear [44].
Text sanitization is a type of text anonymization technique that involves i) detecting and ii) protecting sensitive text spans. Most approaches in the existing literature focus on the detection step, and employ straightforward methods to protect the detected text spans, such as suppressing them or replacing them by generic semantic categories indicating named entity types such as PERSON or LOCATION [21, 46, 47, 18, 25, 9, 30]. However, this relatively crude level of protection can undermine the usability of the protected data for research and analysis, which is the ultimate goal of data sharing. As highlighted in the data privacy literature [19], data protection should aim at balancing the trade-off between the attained level of privacy and the preservation of data utility. For textual data, this means designing effective text sanitization mechanisms that preserve the meaning of the original text as much as possible while fulfilling the privacy needs of the data release.
1.1 Contributions and plan
Large Language Models (LLMs) are particularly effective for text rewriting tasks, among others, due to their ability to understand both the detailed and broader contextual aspects of language. This enables them to rephrase content while preserving its intended meaning, making them valuable for applications like text simplification [22] and editing [41]. Although LLMs can be adapted for text rewriting through fine-tuning on task-specific data, a simpler approach is to rely on instruction-tuned LLMs to directly prompt the model with rewriting instructions.
In this paper, we propose a novel text sanitization approach, named INference-guided Truthful sAnitization for Clear Text (INTACT). INTACT leverages instruction-tuned open-weights LLMs to replace personal information with less specific, but truth-preserving generalizations that retain the original semantic content to the greatest extent possible, while preventing inference attacks from recovering the initial data.
Unlike previous privacy-oriented text rewriting methods [5, 13], our method adopts a two-stage strategy. Given a text span to replace, we first (1) generate a sorted list of generalization candidates at different abstraction levels, and then (2) select the most suitable replacement among those alternatives. Crucially, this selection of the best replacement is done by testing whether the LLM can guess the original span from the replacement and the rest of the text (inference attack). As LLMs have been shown to be effective adversaries in anonymization tasks in previous studies [31], leveraging their predictions is a crucial step towards ensuring the privacy of the generated replacements, particularly truthful ones. Separating these steps not only allows balancing and customizing the privacy-utility trade-off, but it also enables a more granular evaluation of the rewriting process.
We evaluate and compare our method against several baselines on the Text Anonymization Benchmark (TAB) [32], a dataset explicitly tailored for the evaluation of text sanitization methods. As text sanitization is by definition a process that involves the processing of (sometimes highly sensitive) personal data, our approach relies on the use of open-weights models, as those can be run locally without necessitating the transfer of data to third-party cloud solutions. Empirical results show that our method achieves the best privacy-utility trade-off, while keeping the protected data truthful to their original content. We also show how privacy and utility can be measured without reference annotations.
The remainder of the paper is structured as follows. Section 2 provides an overview of the text anonymization task, and discusses related works on text sanitization and text anonymization evaluation. Section 3 presents our text sanitization method, INTACT, and Section 4 describes our utility and privacy evaluation metrics. Section 5 gives implementation details and reports results and evaluation figures. Section 6 provides the conclusions and outlines future research directions.
2 Background and related works
Text anonymization aims to remove or mask any personally identifiable information (PII) that may lead to re-identifying the subject(s) referenced in the text and/or disclose their confidential information.
For detecting PII, the vast majority of automatic text anonymization approaches rely on Named Entity Recognition (NER) [21, 46, 47, 18, 25, 9], which assume that named entities (NEs), due to their specificity, may facilitate re-identification and should therefore be subjected to protection. However, because not all PIIs fit within the limited set of NE types supported by NER tools, NER-based methods may miss a significant amount of disclosive terms [24]. To mitigate this issue, other approaches grounded on the broader notion of disclosure risk have been proposed [17, 30, 35].
The above approaches provide solutions for the detection of disclosive terms, but do not specify how those terms ought to be edited in the document. Two editing strategies are possible. Whereas document redaction directly suppresses or blacks out terms, thereby maximizing privacy at the cost of the utility and readability of the protected documents, text sanitization replaces terms with broader, more abstract generalizations [8]. The latter achieves a better privacy-utility trade-off while preserving the truthfulness of the information, as generalizations, by definition, should retain the core semantic content of the original terms.
The majority of current text anonymization methods use either strict redaction or replace the disclosive terms by their corresponding NE category (such as PERSON or LOCATION), which is a basic form of sanitization [24]. Both strategies retain little to no meaningful information from the original terms. More sophisticated approaches leverage structured knowledge bases, such as ontologies, or, more recently, language models, to retrieve potential replacements. We survey those methods in the following subsections.
2.1 Sanitization from Knowledge Bases
Automatic text sanitization methods have traditionally leveraged structured knowledge bases [10, 3, 2, 37, 38, 36, 17, 29], such as WordNet, the now defunct Open Directory Project (ODP) or YAGO, to retrieve term generalizations to be used as replacements.
They propose different mechanisms to select the most appropriate generalization from a list of candidates. In [17], the authors leverage the Word2Vec [28] word embedding model to identify terms that are semantically related to the individual to be protected. If a term’s embedding has a cosine similarity with the individual’s name above a user-defined threshold, it is considered disclosive, and it is replaced by its most specific generalization (retrieved from YAGO) whose cosine similarity falls below the threshold. In [37, 36] the authors employ a similar approach, but this time selecting the most appropriate replacement through measures of the term informativeness and mutual information. [3] propose to select generalizations so that they encompass a number of specializations (or hyponyms) that would result in $t$ or more plausible documents. [2] employs a contingency table that quantifies the degree of correlation between each pair of textual terms and a taxonomy modeling term generalizations. Finally, [29] rely on human judgments to select the most appropriate generalization for each sensitive term based on its context.
Nonetheless, sanitization methods based on “static” knowledge bases have limited applicability due to their relatively narrow coverage of disclosive terms. This limited coverage is particularly salient for named entities, which are rarely present in such knowledge bases, while they are often one of the most crucial type of terms to sanitize. Moreover, the largest knowledge bases, such as YAGO or ODP may not provide consistent generalizations, as they originate from arbitrary categories, labels and descriptions (such as Wikipedia categories) that do not fulfill the subsumption property; and this may hamper the truthfulness of the sanitized outcomes.
2.2 Sanitization with Language Models
More recently, a number of studies have also explored LLMs for text anonymization, both for detecting and for protecting sensitive text spans. Such models successfully overcome the aforementioned coverage issues of knowledge bases as they can (1) generate replacements for any input span and (2) can ensure those replacements can be inserted in the document in a grammatically meaningful manner.
The MAPA anonymization toolkit [5] is an open source tool capable of both detecting a variety of PII types and replacing them with alternatives sampled from the most likely next tokens based on BERT language models fine-tuned for a variety of European languages and domains including administration, health and law. [1] relies on the same type of models, but without fine-tuning for PII replacement. Those two approaches, however, generate replacements that are restricted to the same token length as the original sensitive term. This restriction has two important shortcomings. The first is that it rules out replacements that are either shorter or longer (in terms of number of tokens) than the original text span. Furthermore, this restriction may be easily exploited by an adversary seeking to infer the original content of the replaced span.
More recently, larger, instruction fine-tuned language models such as those offered by ChatGPT [45] have also been proposed for protecting sensitive information. Instead of fine-tuning, a common way to get these models to perform a task is via an instruction (prompt) and optional examples. Such prompting methods are also integrated in existing de-identification tools, such as Microsoft Presidio. Given a text, the PII detected with the Presidio pipeline can be replaced with GPT 3.5-generated alternatives in a single step [^2] https://github.com/microsoft/presidio/blob/main/docs/samples/python/synth_data_with_openai.ipynb.
None of the LLM-based works mentioned above focus explicitly on generating truthful replacements – that is, replacements that subsume the semantic content of the original text span. As a result, the selected replacement may significantly distort the original terms’ meaning. A study in this direction on a closely related tasks is described in [13], who look at abstracting text spans containing self-disclosure, “the communication of personal information to others”. The categories of personal information employed in this study are a fine-grained version of those typically considered by previous anonymization approaches, and mostly fall under demographic attributes. The sensitive text spans in [13], however, are often phrases and clauses (written in first person) rather than words or named entities. Their proposed task can be thus viewed as a specialized case of protecting sensitive information. PII protection is performed step-wise, where self-disclosure spans are first detected with a BERT model fine-tuned on a manually annotated dataset. The text spans expressing personal information are then re-written with an open-weights instruct LLM (Llama-2-7B) fine-tuned on a dataset with replacements automatically generated using GPT-3.5 and GPT-4.
In contrast, our work leverages open-weights instruct LLMs without additional fine-tuning to iteratively select truthful generalizations from a set of alternatives with different abstraction levels, using inference attacks to test whether a given replacement effectively conceals the initial content of the span.
2.3 Evaluation
A standard evaluation strategy for text anonymization methods is to contrast their outcomes against manual annotations, which are considered as ground truths. The anonymization quality is quantified via the precision and recall metrics, which are used as proxies of the preserved utility and the attained privacy protection, respectively [24, 32, 35, 16, 21].
Precision measures the percentage of text spans replaced by the automatic method that human annotators also deemed disclosive. Because any additional replaced text spans are deemed unnecessary, lower precision figures are interpreted as poorer utility preservation. On the other hand, recall quantifies the percentage of disclosive text spans, as identified by human annotators, that the automatic method replaced. Because untouched disclosive spans are considered privacy failures, lower recall figures are understood as weaker privacy protection.
However, these evaluation metrics are only applicable to assess the quality of the detection step, and are unable to assess the suitability of the replacements used for sanitization. Indeed, evaluating sanitized texts against a manually constructed ground truth is challenging because i) there can be multiple valid (i.e., non-disclosive) combinations of replacements for a given text and, ii) replacement candidates are, quite often, unbounded. Establishing a single ground truth for the evaluation is thus impossible. Moreover, requesting human annotators to provide privacy-preserving replacements (or, at the very least, chose appropriate replacements for predefined lists of candidates) requires even more effort than the –already quite costly– detection of disclosive text spans [29]. As a result, the few sanitization strategies proposed in the literature are largely untested [10, 20, 2, 37, 38, 36, 17, 29].
3 Text Sanitization Method
Our text sanitization approach, INTACT, is based on instruction-fine-tuned LLMs and breaks down the task into two separate steps:
  • 1.
    Generating replacement candidates for disclosive terms with different levels of generalization
  • 2.
    Selecting the most appropriate replacement among those candidates through the use of inference attacks.
The purpose of this two-stage process is to select replacements that satisfy three conditions:
  1. 1.
    they should be privacy-preserving, i.e. they should conceal the original content of the span, also to adversaries seeking to infer their values based on the selected replacements and the context they appear in.
  2. 2.
    they should be truthful, i.e. they should subsume the semantic content of the original span.
  3. 3.
    they should be as informative as possible (to maximize the data utility) given that the two above conditions are satisfied.
A high-level visual overview of INTACT is presented in Figure 1 together with an example replacement generation and selection outcome.
Figure 1: Overview of the text sanitization process with INTACT, together with a sample system output.
We will assume throughout this paper a document $D$ in which we have detected a set of text spans $T=\{t_{1},t_{2},\allowbreak\ldots,t_{n}\}$ that may disclose personal information and ought to be edited. The task of detecting those PII-related spans in text documents is outside the scope of this paper, but has already been studied in numerous works, such as [27, 24, 32]. Common approaches to this task rely on standard NER tools or sequence labeling models tailored for detecting PII-related information. We will also assume that each of those detected text spans $t_{i}$ in $D$ has been tagged with a corresponding entity type label $l_{i}$, such as ORG or PERSON.
3.1 Replacement Generation
In this first step, we prompt an LLM to generate a list of generalized replacement candidates for the provided text span in its context. We explicitly request replacements with different levels of abstraction that the LLM is then required to sort from most to least specific. This sorting is done within the same step in order to limit computational costs. To enhance the quality of the generated candidates, we also include a one-shot example containing a sentence with a target span (of the same label as the span to sanitize) and some possible generalizations. We provide the text span in its local context, defined here as the sentence it occurs in.
More formally, given an open-weights instruction fine-tuned language model $\texttt{LLM}_{\texttt{instr}}$ and an input text span $t_{i}$ occurring in $D$ that should be sanitized, we prompt the LLM to generate a list of $m$ replacements candidates $C_{i}=\{c_{i,1},c_{i,2},\ldots,c_{i,m}\}$, where the candidates $c_{i,j}$ are sorted by order of abstraction (from the most specific to the most generic). The list of replacement candidates $C_{i}$ for the text span $t_{i}$ is thus defined as:
$C_{i}=\texttt{LLM}_{\texttt{instr}}(p_{C},e_{C},t_{i},s_{i})$ (1)
where $p_{C}$ is the prompt used to instruct $\texttt{LLM}_{\texttt{instr}}$ to generate the candidates, $e_{C}$ is the one-shot example, $t_{i}$ is the input text span that should be edited, and $s_{i}$ is the sentence in $D$ containing that text span.
The text spans $\{t_{1},t_{2},\allowbreak\ldots,t_{n}\}$ are processed according to their order of occurrence in $D$. For entities that occur multiple times through the same document $D$, we only generate replacements for the first mention of that entity and reuse the selected replacement for all subsequent mentions.
For direct identifiers – such as person names or phone numbers – which can reveal identity even in isolation, replacement candidates are not obtained through Eq. 1. Instead, in line with how direct identifiers are handled in statistical disclosure control (SDC) [19], $C_{i}$ of direct identifiers are assigned to either:
  • 1.
    a generic token (e.g., “***”);
  • 2.
    an entity type (e.g., PERSON) corresponding to the label $l_{i}$;
  • 3.
    a more human-readable variant of this category (e.g., “the person”).
This not only helps maintaining privacy, but it also prevents accidental identification, i.e., generated names or codes matching actual entities.
3.2 Replacement Selection
The second step of the method identifies the best replacement $c_{i}^{*}$ among the possible replacement candidates $C_{i}$ for the original text span $t_{i}$. The best replacement is here defined as the most specific candidate that does not reveal, directly or indirectly, the original content of the span $t_{i}$. In other words, the best replacement should be the most informative replacement among those that do not constitute a disclosure risk. To determine whether a replacement candidate $c_{i,j}$ constitutes a disclosure risk, we carry out an inference attack – that is, we use the instruction-tuned LLM to produce a set of guesses for the original span given some context. These guesses are then compared to the original span to assess whether the replacement is acceptable (i.e. if none of the guesses match the original span) or whether it would allow an adversary to find back the original text span. Crucially, those guesses are generated by taking into account not only the replacement candidate, but the full (edited) document as context, in order to allow the LLM to take advantage of all information expressed in the document to predict the original content of a given span.
Inference attack
For each text span $t_{i}$ and replacement candidate $c_{i,j}$ for this span, we generate a set of $p$ guesses $G_{i,j}=\{g_{i,j}^{1},g_{i,j}^{2},\allowbreak\ldots,g_{i,j}^{p}\}$ using the instruction-tuned LLM:
$G_{i,j}=\texttt{LLM}_{\texttt{instr}}(p_{G},e_{G},c_{i,j},D^{\prime})$ (2)
in which:
  • 1.
    $p_{G}$ is the prompt employed for the generation of guesses.
  • 2.
    $e_{G}$ is an one-shot example composed of a short purpose-written text, with one span substituted with a synthetic LLM replacement and a set of synthetic LLM guesses.
  • 3.
    $c_{i,j}$ is the candidate replacement for the original input text span $t_{i}$.
  • 4.
    $D^{\prime}$ is the current version of the sanitized document. More precisely, $D^{\prime}$ is obtained from the original document $D$ by replacing each original text span $t_{j}$ with its best replacement $c_{i}^{*}$, if it has already been selected, or with the most specific replacement candidate $c_{i,1}$ if is has not.
Matching
We then need to define what constitutes a “successful” guess from the point of view of the adversary. A naive approach would be to define a guess as successful if it returns the exact same string as the original text span. This approach, however, fails to account for guesses that are closely related but not entirely identical to the original string. To capture those variations, we define a $\small\textsc{match}(t_{i},g_{i,j}^{k})$ condition, which returns true if any of the following conditions hold:
  • 1.
    Lemma Overlap: There is an overlap between the set of lemmas (base forms of words, e.g., dog for dogs) of $t_{i}$ and $g_{i,j}^{k}$.
  • 2.
    N-gram Overlap: If $t_{i}$ is a named entity, an overlap exists between the n-grams (sequences of $n$ characters) of $t_{i}$ and $g_{i,j}^{k}$.
Non-alphabetic tokens, stopwords and high-frequency lemmas being sanitized are excluded from both $t_{i}$ and $g_{i,j}^{k}$ when computing lemma overlap. Acronyms are handled by merging the first letters of all title-cased words in $t_{i}$ into one word, doing the same for $g_{i,j}^{k}$, and expanding the respective lemma lists with these automatically coined (potential) acronyms. In all comparisons, the text spans are normalized to lowercase characters.
Based on this matching function, we can then define a replacement candidate $c_{i,j}$ as risky from a privacy perspective if at least one of the guesses matches the original text span $t_{i}$. More formally, we can thus define the following boolean function:
[…]
3.3 Full sanitization process
The selection of the best replacement $c_{i}^{*}$ among possible alternatives proceeds as follows:
  • 1.
    For a particular text span $t_{i}$ to sanitize, we generate the replacement candidates $C_{i}$ according to Eq. (1)
  • 2.
    For each replacement candidate $c_{i,j}\in C_{i,j}$, we generate the guesses on the content of the original text span based on Eq. (2)
  • 3.
    The candidates $c_{i,j}\in C_{i}$ are then examined in the order they were generated by the LLM. For each candidate, we compare its corresponding guesses to the original text span $t_{i}$.
  • 4.
    Finally, we select as replacement $c_{i}^{*}$ the first replacement candidate $c_{i,j}$ for which risky_replace returns false. If no candidate is found, as a fallback strategy, the text span is replaced by the name of its corresponding entity type $l_{i}$, such as LOCATION [^3] In the absence of entity labels, a placeholder such as “***” can be used..
This process is detailed in Algorithm 1.
Algorithm 1 Text sanitization procedure sanitize(D).
An input text document $D$
A sanitized document $D^{\prime}$
An instruction-tuned $\texttt{LLM}_{\texttt{instr}}$
# Detect the disclosive text spans in $D$ and their semantic categories
# (the implementation of this function is outside the scope of this paper)
[…]
# Step 1: find replacement candidates for each disclosive span
for $t_{i}\in T$ do
     Extract sentence $s_{i}$ in $D$ in which $t_{i}$ occurs
     # Generate replacement candidates with prompt $p_{C}$ and one-shot $e_{C}$
     $C_{i}\leftarrow\texttt{LLM}_{\texttt{instr}}(p_{C},e_{C},t_{i},s_{i})$
end for
Initialize $D^{\prime}\leftarrow D$ where all spans $t_{i}$ (for $1\leq i\leq n$) are initially replaced by their most specific replacement candidate $c_{i,1}$
# Step 2: conduct inference attacks to determine best replacement
for $i\leftarrow 1\text{ to }n$ do
     for $j\leftarrow 1\text{ to }m$ do
         # Generate guesses about the original span based on candidate $c_{i,j}$
         $G_{i,j}=\texttt{LLM}_{\texttt{instr}}(p_{G},e_{G},c_{i,j},D^{\prime})$
     end for
     # Select candidate that is most specific and not risky
[…]
     # Fall back to a default value if all replacements were risky
[…]
         $c_{i}^{*}\leftarrow$ semantic label $l_{i}$ or placeholder
     end if
     Update $D^{\prime}$ by replacing $t_{i}$ by $c_{i}^{*}$
end for
return $D^{\prime}$
4 Evaluation metrics
As discussed in Section 2.3, the current metrics for text anonymization evaluation are not suitable to assess the quality of sanitized replacements. To address this shortcoming, we propose two novel automatic evaluation metrics. Instead of relying on manual annotations, our metrics quantify data utility as a function of the proportion of semantic content preserved in the selected replacements relative to the original text span, and privacy protection as the inverse of the observed residual re-identification risk.
4.1 Utility assessment
The utility purpose of anonymized data is to produce outcomes in downstream tasks that closely align with those resulting from the original data.
In SDC for structured databases [19], this alignment is evaluated by measuring the similarity between the anonymized and the original data, as the two are known to be correlated. However, while quantifying similarity for structured numerical data is straightforward, doing so for unstructured text data is more challenging.
We propose a new utility metric for text anonymization that quantifies the similarity between an original document and its sanitized version from a semantic perspective, as the preservation of the semantic content of the text is directly related to its utility. Our proposal assesses the impact on utility of these replacements based on two factors:
  1. 1.
    the information content of the disclosive text span to be protected within the original document, which accounts for the amount of semantics encompassed by the original text span.
  2. 2.
    the semantic similarity of the replacement relative to the original text span, which expresses the proportion of semantic content preserved by the replacement.
4.1.1 Information content
In natural language processing, the informativeness of a text span, or its information content (IC), can be expressed according to Shannon’s information theory as the negative logarithm of its probability of occurrence[40]:
$\small\textsc{IC}(t)=-log(\Pr(t))$ (3)
This implies that the less probable (or predictable) a text span $t$ is, the more information it provides. Conversely, if the text span is highly probable (or predictable), its information content is low.
Under this premise, we calculate the information content of a document $D$ (that is, the amount of semantics it conveys) by summing the IC values of all its text spans. We name this text information content (TIC):
$\small\textsc{TIC}(D)=\sum_{t\in D}{\small\textsc{IC}(t)}$ (4)
As all text spans we consider the set of non-overlapping segments that are either a noun phrase, a named entity or a word other than a stopword.
We then measure the relative information content (RIC) of a text span $t$ within a document $D$ as the proportion of the text span’s IC relatively to the TIC of $D$:
$\small\textsc{RIC}(t,D)=\frac{\small\textsc{IC}(t)}{\small\textsc{TIC}(D)}$ (5)
Various methods are possible to practical compute the probability $\Pr(t)$ in Eq. 3, such as using their frequency of appearance within a corpus [34] (even leveraging the Web as corpus [37, 35]), considering the number of hypernyms and/or hyponyms within an ontology [39, 7], or evaluating their predictability with a language model[32]. The latter approach utilizes the general language understanding capabilities of language models (specifically BERT [12]) to predict words based on context. The model receives the context of the text span (e.g., the surrounding sentence) with the target text span substituted by the [MASK] token. The probability of the text span is then determined by the model’s probability of predicting it from the context. For multi-token text spans, the minimum probability is selected. This predictability technique is the only estimate of $\Pr(t)$ that accounts for compositionality, which states that the meaning of an expression is derived from the meanings of its parts and the rules for combining them. Moreover, this technique is also the only capable of estimating probabilities for any text span, not requiring it to be present in a specific corpus or ontology.
As in [32], we rely on BERT (bert-base-uncased [^4] https://huggingface.co/google-bert/bert-base-uncased) for computing the IC of our text spans. However, contrary to [32], here we must compute the IC for all text spans within a document. Using the method proposed in [32] would involve substituting all text spans with [MASK] simultaneously, depriving the model of the necessary context to make accurate probability predictions. To solve this issue, we alternate the [MASK] substitution across text spans. Specifically, placing [MASK] for one of every $N$ text spans, using the remaining spans as context. This technique requires making multiple predictions, each time shifting the [MASK] placements. For instance, with $N=2$, two predictions are needed: one with [MASK] applied to the odd text spans while using the even spans as context, and another with [MASK] applied to the even spans while using the odd spans as context. Increasing $N$ linearly increments the number of predictions (and thus the computational cost), but it is expected to enhance the quality of the probability estimates by providing more context for each prediction.
4.1.2 Semantic similarity
As introduced above, the semantic similarity between an original text span and its replacement indicates how well the semantic content of the former has been preserved by the latter. A well-established measure of semantic similarity between text spans consists of calculating the cosine similarity between their embeddings in a language model [43, 11, 4].
A cosine similarity of 1 corresponds to identical spans, while 0 indicates no particular similarity, and -1 depicts completely unrelated text spans. Given that replacements are not expected to be unrelated to the original text span, we constrain the similarity calculation to the $[0,1]$ range:
$\small\textsc{Similarity}(t_{i},c_{i}^{*})=\max(0,\small\textsc{cos\_sim}(\small\textsc{embed}(t_{i}),\small\textsc{embed}(c_{i}^{*})))$ (6)
Where $t_{i}$ is the original text span, $c_{i}^{*}$ the selected replacement span, and embed is the embedding model.
As embedding model, we leveraged paraphrase-albert-base-v2555https://huggingface.co/sentence-transformers/paraphrase-albert-base-v2 in our experiments, an off-the-shelf language model trained with Sentence-BERT [33] that has been shown to achieve state-of-the-art results in multiple benchmarks for semantic similarity [33, 23].
4.1.3 Text Preserved Similarity
Building upon the aforementioned measures, we now introduce text preserved similarity (TPS), an utility metric for text sanitization:
$\small\textsc{TPS}(D,D^{\prime})=\sum_{(t_{i},c_{i}^{*})\ \in\ (D,D^{\prime})}\small\textsc{RIC}(t_{i},D)*\small\textsc{Similarity}(t_{i},c_{i}^{*})$ (7)
Where $D^{\prime}$ represents the sanitized document, $D$ the original document, $t_{i}$ the original span and $c_{i}^{*}$ its selected replacement.
For unedited text spans, $c_{i}^{*}$ is identical to $t_{i}$, resulting in a similarity of 1 and thus no loss in utility. TPS can be interpreted as the arithmetic mean of semantic similarities for all text spans, weighted by their relative informativeness. The TPS metric range from 0, indicating a loss of all original content, to 1, where $D^{\prime}$ conveys precisely the same meaning as $D$ (e.g., no text spans have been altered).
4.2 Privacy assessment
The ultimate goal of data anonymization is to minimize the re-identification risk. Considering that the greatest disclosure risk comes from the combination of disclosive text spans [24], a sound privacy evaluation should consider the residual re-identification risk of the chosen replacements as a whole. In [26] we proposed a privacy metric for anonymized text, named text re-identification risk (TRIR) to precisely account for this.
More specifically, TRIR involves simulating a plausible re-identification attack on the protected documents by leveraging, as background knowledge, a collection of public documents of a population of individuals that encompasses the actual individuals to be protected. For example, if the individuals to be protected are hospital patients and the anonymized documents are their medical reports, all the citizens of the hospital’s town could be the population, and the public documents leveraged as background knowledge could be a collection social media posts of those citizens.
To construct the attack, a language model with a classifier head is first trained to exploit this background knowledge to predict the individual corresponding a given document. This classification model is thus trained to identify the information most closely linked to each individual from the public documents. When presented with an anonymized document, the model can then leverage the learned knowledge to re-identify the corresponding individual. If the predicted public individual is in fact the corresponding individual to be protected, re-identification is successful. TRIR is therefore computed as the re-identification accuracy on the set of anonymized documents.
5 Experimental results
How does the approach to replacement generation and selection outlined in Section 3 compare to previous methods? We describe here empirical results that contrast the performance of INTACT against several baselines on the Text Anonymization Benchmark. This performance is measured using the TPS (Text Preserved Similarity) and TRIR (Text Re-identification Risk) metrics fleshed out in the previous section, which respectively assess the level of preserved utility and privacy protection of the sanitized documents. Furthermore, we also conduct a manual evaluation to provide a finer-grained assessment of the truthfulness and the abstraction level of the selected replacements [^6] All code and data employed in the experiments are available at https://github.com/IldikoPilan/text_sanitization..
5.1 Dataset
We employed the documents and manual annotations of the Text Anonymization Benchmark, which encompasses 1,268 English documents extracted from court cases of the European Court of Human Rights (ECHR). The Procedure section of these documents, which is rich in personally identifying information, was manually annotated by 12 law students, who tagged disclosive text spans and specified their corresponding entity type by following privacy-oriented annotation guidelines [32].
On average, 11% of the text spans in the documents were considered as PII by the annotators. The entity types of the corpus are: CODE, (organization), DATETIME, LOC (location), QUANTITY, PERSON, DEM (demographics) and MISC (any personal information not falling in the previous categories). As the proposed approach does not require any fine-tuning, we conduct our experiments on the full set of documents from the TAB corpus. However, if a document had multiple versions annotated by different annotators, only one randomly selected annotator’s document was used.
5.2 Implementation details
We describe below the technical details involved in the application of INTACT to the TAB corpus.
Replacement generation
For all semantic categories other than CODE and PERSON, we generate replacement candidates (see Eq. (1)) with a 4-bit quantized Mistral 7B Instruct v.0.2777https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 model with multinomial sampling, a temperature of 0.3 and with a limit of 512 new tokens. The number $m$ of replacements is set to five. The document is segmented into sentences using Spacy [^8] https://spacy.io/. Each target PII span $t_{i}$ is delimited by double square-brackets in the sentence. The complete instruction formatted as a multi-turn chat and the example sentences used are presented in Tables 3 and 2 in the Appendix.
Heuristic rules
For the CODE and PERSON categories, which are direct identifiers, the replacement consists of the entity type label and a running number (e.g., CODE_0) thereby maintaining reference links between different mentions of the same underlying entity.
We also implement rule-based generalizations for DATE for some common date formats and use LLM-based replacements only as fallback strategy. This is motivated by the well-known weakness of many current smaller-sized LLMs to handle numeric information and the potentially identifying nature of e.g., birth dates. The starting level of abstraction for the list of replacements for dates reflects the specificity of the original date. For example, for a date consisting of day+month+year, the most specific generalizations is month+year, but for an original date span of month+year, the most specific replacement is season+year.
Replacement selection
For selecting replacements, we employ the same LLM used for generating the candidates (the aforementioned Mistral 7B instruct model). The number of guesses is set to five and the same delimiters for the span being processed are used as for generation, see the prompt in Table 4 in A.2. When all replacement candidates are guessed, the entity type and a running number (e.g., ORG_2) are used as fallback.
Matching function
To assess whether the LLM was able to correctly guess the original text span based on its replacement and the rest of the document context, we defined a matching function (see Section 3.2). This function first checks the existence of a lemma overlap (either exact or partial) between the guess and the actual span. Lemma overlap for DATETIME entities is handled somewhat differently, as generalization here is mostly rule-based and results in re-using a subset of the original information (e.g.August 2003 for 3 August 2003). We therefore prompt the LLM to always try to guess exact dates and we only consider a DATETIME span guessed if all lemmas between the original span and a guessed span are the same. Allowing a partial overlap between the original and the guessed lemmas might be an acceptable type and level of generalization also in some other cases involving other entity types (e.g., guessing only the country when the original span specifies both the city and the country). We also check for n-gram overlap if the text span corresponds to a named entity. We choose 4-grams to account for words with the same, longer base (stem), e.g., Turkey - Turkish.
5.3 Baselines
We compare INTACT with four alternative replacement strategies :
  1. 1.
    Suppression: complete removal of the disclosive text span.
  2. 2.
    Entity replacement: replaces the text span by its corresponding named entity type (e.g., PERSON, CODE, ORG…).
  3. 3.
    Presidio’s synthetic replacement [^9] https://github.com/microsoft/presidio/blob/main/docs/samples/python/synth_data_with_openai.ipynb: replaces the text span entity type with a synthetic alternative (e.g., “[LOCATION]$\rightarrow$“Spain”) generated with GPT-3.5-turbo As mentioned in Section 2.2, because replacements are not intended to be truthful, they are expected to notoriously distort the document’s meaning.
  4. 4.
    Dou et al. replacement: replaces the text span by one of the three alternative replacements suggested by the model from [13]. [^10] For each text span, one of the three candidates is randomly chosen (as the three suggestions of this model are not ordered). The exact model can be accessed at https://huggingface.co/douy/Llama-2-7B-lora-instruction-ft-abstraction-three-span.
We also evaluate the performance of two variants of the INTACT approach, where we retain the replacement generation approach of Section 3.1, but rely on an alternative method for selecting the final replacement among possible candidates. Those two variants are:
  • 1.
    Least-specific replacement: substitutes the text span with the last/more general candidate from the sorted list of replacement candidates.
  • 2.
    Most-specific replacement: substitutes the text span with the first/more concrete candidate from the sorted list of replacement candidates.
5.4 General statistics
Figure 2 shows the frequency of which candidate was finally selected with the INTACT approach from the list of five replacement candidates on the TAB corpus. The candidates are sorted by the LLM by order of abstraction, from the most specific to the most generic/abstract. In a little over half of the cases (53%), the most specific (first) replacement candidate was chosen [^11] For DATETIME entities, the frequent selection of Candidate 1 and 2 can be attributed, in part, to the stricter definition of what constitutes a ”successful” guess, where only complete lemma overlap was considered as guessed..
In about 22% of text spans in the TAB corpus, all five replacement candidates were guessed by the LLM and the entity label was thus used to replace the text span. This default replacement strategy was particularly common for entities labeled as ORG and MISC, amounting to almost half of text spans ending up in this default case, as detailed in Table 5.
Figure 2: Frequency of INTACT’s replacement choice for TAB. Entity Type is selected when all five candidates are guessed by the inference attacks, leading to a default strategy of replacing the text span by its entity type label (e.g. PERSON).
5.5 Utility evaluation
Intrinsic Evaluation
For evaluating utility preservation, we first apply the TPS metric proposed in Section 4.1 between the original and protected documents. [^12] To determine the value for the hyper-parameter $N$ specifying the spacing between text spans replaced by [MASK] tokens (see Section 4.1.1), we conducted tests across multiple values (i.e., $N\in[2,4,6,8]$), and took $N=6$ as a good trade-off between runtime and the considered context size. Figure 3 depicts the TPS values obtained for the different strategies. Since all of them protected the same text spans (that is, those manually annotated in TAB), discrepancies in results arise from the varying semantic similarities between the original text spans and their replacements.
Figure 3: TPS for the different replacement strategies
The worst utility corresponds to the suppression strategy, whereas utility figures strictly increase as we employ more specific/utility-preserving replacements, that is: entity types $\rightarrow$ Presidio replacements $\rightarrow$ least-specific replacements $\rightarrow$ Dou et al. replacements $\rightarrow$ INTACT replacements $\rightarrow$ most-specific replacements. As a matter of fact, the average semantic similarities for the different replacements were 0.33, 0.38, 0.49, 0.49, 0.51, and 0.62, respectively.
As mentioned in Section 5.1, about 11% of the text spans in TAB were marked by the annotators as expressing direct or indirect PII. The suppression of those PII-related text spans causes a 22% drop in TPS, suggesting these spans are, on average, twice as informative as the non-protected spans. As illustrated in Figure 2, the two most common replacement candidates selected by INTACT are respectively the first candidate of the sorted list of replacements or the default entity type, which explains why its TPS falls between that of entity replacements (i.e., 85) and the most-specific replacements (i.e., 92).
Extrinsic Evaluation
To further validate the results, we also evaluated the utility preserved by the different strategies on a downstream task: document clustering. To quantify the task-oriented utility preservation, we compared the clusters obtained from the original documents (which we considered the ground truth), with those resulting from the protected documents under the different strategies. Utility preservation was then computed as the similarity between both clusterings, as similar clusters suggest that equally similar conclusions can be reached from the analysis of the protected documents.
To ensure that clustering leverages the semantic content of the document, we followed the method from [42], which relies on BERT [^13] More specifically the bert-base-cased model, see https://huggingface.co/google-bert/bert-base-cased. [12] to generate document embeddings by averaging the embeddings over all individual tokens. Once the document embeddings were obtained, they were clustered using the K-means++ algorithm [6] with $K=4$. The value of $K$ was selected based on the optimal inertia achieved when clustering the original documents. To mitigate the effect of the random initialization, the clustering for each document set was repeated 50 times, choosing the results with the lowest inertia. To quantify the similarity between clustering results, we used the Normalized Mutual Information (NMI) [14]. For NMI calculation, the average of 5 runs of the aforementioned clustering process was computed. In summary, each clustering was performed 250 times ($50\times 5$) to minimize the impact of the random initialization.
Clustering-based utility evaluation for the considered strategies are shown in Figure 4. If we compare the results of this clustering-based utility evaluation with TPS, we obtain a strong Pearson correlation $r=0.864$ ($p$-$value=0.012$). This shows the effectiveness of TPS as an a priori estimator of the analytical utility of the protected documents in downstream tasks.
Figure 4: NMI for the clustering resulting from the evaluated replacement strategies
Method-level differences offer additional insights into how various approaches impact data utility. For instance, as one can observe from Figure 4, entity replacements obtained a higher NMI than least-specific replacements, while the opposite holds for TPS. Similarly, INTACT achieves a NMI higher to that of the most-specific replacements, yet their TPS is lower. A common factor in both cases is the use of entity types. Although generated replacements are more similar to the original text spans when examined individually (i.e., higher TPS), some tend to distort the general document content more than entity types (i.e., lower NMI). Such distortion likely arises when a generalization leads to overly broad replacements that disrupt the linguistic register and structure of the document (as exemplified by the three least-specific replacements in Example 1), or even unfaithfully abstract its initial content. The strategy of using entity types when all generated replacements are guessed (which is frequent, as shown in Figure 2) in fact resulted in a NMI that surpassed that of the most-specific replacements. It seems that, when the language model fails to generate replacements that sufficiently protect privacy, it also struggles to produce options that effectively preserve utility.
Discrepancy between TPS and NMI is significantly more pronounced with Dou et al. and least-specific replacements, which achieved a TPS nearly identical to INTACT but resulted in a much lower NMI. The replacements for these methods are observed to often be lengthy sentences that, while resembling the original span (i.e., high TPS), tend to disrupt more the overall text meaning (i.e., low NMI).
Presidio’s synthetic replacements achieved a notably high NMI, outperforming both entity-based and least-specific replacements, likely due to their better integration within the text. Nonetheless, even when using a more advanced model than the one used to generate our replacements (i.e., GPT-3.5-turbo $>$ Mistral-7B [^14] https://lmarena.ai/?leaderboard), its non-truthful replacements still yielded a lower NMI compared to our most-specific and INTACT’s selected replacements.
5.6 Privacy evaluation
We evaluated the residual disclosure risk of the strategies by using the TRIR metric introduced in Section 4.2161616For TRIR calculation, we leveraged the ready-to-use implementation available in the corresponding GitHub repository  [^15] https://github.com/BenetManzanaresSalor/TextRe-Identification with the default attack hyperparameters..
The use of TRIR requires the specification of a body of background knowledge on the basis of which the re-identification risk is derived. We relied on a corpus of 13,759 rulings published on the ECHR website to define this background knowledge. The annotated court decisions from TAB were derived from the same source, this corpus constitutes therefore a strict superset of the TAB documents. However, assuming that each adversary has access to this full corpus of rulings is somewhat unrealistic, as such an adversary would already have access to all original documents, and would therefore have no need to re-identify the sanitized versions of those same documents. On the other hand, considering that the knowledge of the adversary strictly encompasses the protected individuals is also unrealistic, as they would rarely known the exact population sample in the protected dataset. Taking this into consideration, we defined the background set as a random superset of TAB encompassing twice the number of individuals/documents (i.e., $1,268\times 2=2,536$), and considering the first 10% of non-Procedure text from the rulings.
The TRIR for each strategy is reported in Figure 5. Notice that, as any ex post privacy metric, TRIR measures re-identification risk rather than actual re-identification. Subsequently, the $\approx 10\%$ figures resulting from all methods do not mean that 10% documents have been unequivocally re-identified, but that attackers have a 10% chance of accurately guessing the protected individual, as they cannot verify the correctness of their predictions.
Figure 5: TRIR for different replacement strategies
As expected, the replacements with the most-specific LLM generalisations yield the worst TRIR scores. The difference between this approach and the suppression strategy is, however, relatively small (2.2 percent points). In other words, the reliance on informative replacements (generalisations) makes it possible to substantially improve utility without substantially impairing privacy with respect to the strictest protection, i.e., suppression. INTACT performed particularly well on this balance, achieving a relatively low TRIR overall – second lowest when considering only generated replacements –, while delivering the second-highest TPS and the highest NMI. Compared to both the least-specific and most-specific replacements, INTACT delivered lower TRIR alongside comparable or superior TPS and NMI.
On the other hand, Presidio’s and entity replacements incurred the lowest risks because of the non-truthful replacements of the former (see Section 5.7), and the abstract replacements of the latter tended to confuse the re-identification model. This reduced the re-identification accuracy even more than suppression, because the re-identification model relies on concepts found in the background knowledge of the individuals targeted for re-identification. While suppression removes part of that exploitable information, Presidio’s synthetic and entity replacements substitute it by new information that is rarely found in any background knowledge. The scenario differs for truthful replacements –like INTACT or those from Dou et al.–, where part of the meaning of the original concept is preserved, and therefore it is possible to connect it with the corresponding individual’s background knowledge, which sightly increases TRIR.
However, regardless of the actual TRIR, using non-truthful replacements as Presidio does, may allow an attacker to differentiate between protected and non-protected terms, thereby providing an advantage that can be exploited to conduct case-specific attacks focused on inferring only the protected terms. Also, an obviously protected text, as it is the case for entity replacements, which use abstract terms, raises awareness of the document’s sensitivity and attracts more attention as a target for attacks [8].
Dou et al. replacements achieved the second-highest TRIR, closely matching that of the least-specific replacements. In fact, this similarity extends consistently across all three metrics –TPS, NMI, and TRIR–, which suggests that this method is similar to selecting the least-specific candidate generated by our approach.
5.7 Truthfulness and abstraction level
An important characteristic of the INTACT approach is its emphasis on producing truth-preserving replacements whose semantic content subsumes the content of the original text span. Such replacements should therefore correspond to generalisations that encompass and abstract away the original text span. However, the truthfulness and abstraction level of the replacements are nuances that are challenging to evaluate automatically in a reliable manner. Their impact in combination is reflected in the results of our automatic metrics, but these do not provide explicit information about replacement truthfulness and abstraction level separately.
To address this, we conducted a manual evaluation focused on the truthfulness and abstraction level of the selected replacements. This evaluation was performed on a subset of TAB documents, and comprised 600 instances, each consisting of a PII span mention with 70 characters preceding and following it as context. For each document, we randomly sample a mention for each entity and then select a random subset of all these sampled mentions, using an equal size of 100 instances per each entity type, excluding the PERSON and CODE categories. The instances were rated by a single evaluator with NLP background, who was also allowed to look up concepts on the internet.
The selected replacements are assessed using two criteria: abstraction level and truthfulness. A replacement is considered more abstract if it constitutes a hypernym or provides fewer details than the original PII span. A replacement is truthful if it retains the same overall content of the original span. When approximate equivalence can be assumed in the context of the span (e.g.28 days and 1 month), it is counted as truth-preserving. Partially true spans (e.g., containing two dates) are not considered truthful.
Table 1 presents the manual evaluation results for Presidio, Dou et al.’s model [13] and INTACT. According to this manual evaluation, 81.33% of the assessed replacements were less specific and 93.17% truthful with INTACT. In contrast, Presidio’s manual evaluation reports 36.16% less specific and just 19.67% truthful replacements. Dou et al.’s abstraction model also produced mostly more abstract and mostly truthful replacements (78.33% and 95.16% respectively). However, replacements that were truthful but not more abstract amounted to 20.16% for Dou et al., which is 5% higher than for INTACT, where the iterative inference attacks are meant to prevent the use of too specific replacements. This supports the TRIR-based automatic evaluation results (Figure 5) which show a higher privacy risk posed by Dou et al. than INTACT.
Table 1: Manual evaluation of replacements on truthfulness and abstraction level for Presidio, Dou et al and INTACT.
Approach Not-truthful Truthful
Same level of More Same level of More
abstraction abstract abstraction abstract
Presidio 61.83% 18.5% 2% 17.67%
Dou et al. 1.5 % 3.33% 20.16% 75%
INTACT 3.5% 3.33% 15.17% 78%
The results from Table 1 show that smaller, open LLMs are well suited to finding generalized replacements in terms of abstraction level and, especially, truthfulness, even without specific fine-tuning for this task. For both INTACT and Dou et al., however, the generation of more abstract replacements was somewhat more challenging to achieve for the DEM and the MISC categories, where text spans are more often specialized or technical terms. The LLMs showed some tendency to produce definitions, explanations or paraphrases for these terms. Sometimes there was still a degree of abstraction between the different alternatives, but in some cases, none of the suggestions were generalizations, e.g., unemployed resulted in the INTACT replacements: Out of work, Jobless, Inactive in the labor force, Economically inactive, Without employment. We also noticed that the truthful and more abstract replacements obtained with Dou et al. were often more generic (e.g., in a specific period for replacing a DATETIME entity) than the ones for INTACT. This type of semantic loss is likely one of the reasons why INTACT outperformed Dou et al. in the clustering-based utility evaluation results presented in Figure 4.
In comparison, Presidio’s replacements were much less consistent, with a vast majority of them being non-truthful and with arbitrary degrees of abstraction. The lack of truthfulness and consistency is expected to significantly distort the document’s meaning. Moreover, we noticed a substitution bias towards the American-related sources, with the nationality of defendants in the TAB corpus changed from Turkish, Polish and British to Brazilian, Canadian and Mexican. Even though these distortions contribute to preventing re-identification (as shown in Section 5.6), it also seriously hampers secondary uses, as it might invalidate conclusions drawn from per-document analysis.
The results from this manual evaluation therefore show that the two-stage approach of INTACT leads to replacements that are truth-preserving in the vast majority of cases, and also typically more abstract than the original text span. As non-truthful replacements are expected to impair utility, while replacements at the same level of abstraction may compromise privacy protection, those results are in line with the ones obtained with the automated evaluation results of the previous sections.
6 Conclusions and future work
We presented INTACT, a two-stage approach to automatically sanitize disclosive text spans in documents by leveraging instruction-tuned large language models. In contrast with related works, our approach is not limited by the narrow scope of structured knowledge bases and focuses on generating truthful replacements, which are of outmost importance in secondary uses of the data, such as medical research, insurance or law. We ensure the replicability of our method by using open-weights LLMs, which can also be run locally to guarantee private and secure handling of sensitive data. Empirical results show that our method provides the best privacy/utility trade-off, while avoiding distorting the original semantic content of the document as other LLM-based approaches do.
We have also proposed automatic utility and privacy metrics for text anonymization, which capture the notion of semantics and disclosure risk better than standard precision and recall metrics, and do not require (costly) manual annotations as ground truth.
In future work, we aim to enhance both the ordering of replacement candidates and their grammatical integration within sentences. Even though a grammatically correct output is specifically requested in the replacement generation prompt, some replacements would still produce ungrammatical phrases (e.g., double prepositions like in in) if inserted, partly due to inconsistencies in the inclusion of these within the manually annotated spans of the dataset used. Additionally, while the replacement suggestions generally range from the most specific to the most abstract, the ordering isn’t always consistent. We also plan to explore additional methods for matching the original spans against those guessed by the LLM. These enhancements could potentially be implemented through additional LLM steps to further improve the quality of the sanitized text.
Acknowledgments
We acknowledge support from the Norwegian Research Council (CLEANUP project (http://cleanup.nr.no/), grant nr. 308904), the Government of Catalonia (ICREA Acadèmia Prize to D. Sánchez, and grant 2021SGR-00115), MCIN/AEI/ 10.13039/501100011033 and “ERDF A way of making Europe” under grant PID2021-123637NB-I00 “CURLING”, and the EU’s NextGenerationEU/PRTR via INCIBE (project “HERMES” and INCIBE-URV cybersecurity chair).
Declaration of generative AI and AI-assisted technologies in the writing process.
During the preparation of this work the author(s) used ChatGPT (with GPT-4o) in order to improve readability and fluency in parts of the text. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
Appendix A LLM-based Text Sanitization
A.1 Replacement Generation
Table 2 shows the synthetic examples used for obtaining the LLM-based replacements per semantic category.
Table 2: Example sentences and replacements used in the one-shot in-context learning setup for replacement generation.
Category Sentence context Sorted replacements
ORG John Smith often volunteered in [[Sunrise Psychiatric Hospital]]. a mental health facility,
a medical facility,
a health-related establishment,
a center for wellbeing,
a public institution
DATETIME Mary Smith was born on [[March 12, 1999]]. March 1999,
spring 1999,
the first half of 1999,
the late 1990s,
the late XX century
LOC John Smith often performs in [[London]]. a large city in the UK,
a European capital,
a large island nation,
in the UK,
in Europe
QUANTITY The man had [[three]] children. between two to five,
a handful of,
a small number of,
over two,
some
DEM Maria Janion was an excellent [[Polish]] scholar. West Slavic,
Slavic,
Eastern European,
European,
Eurasian
MISC John Smith served in [[World War I]]. a military conflict in the first half of the 1900s,
a military conflict in the 20th century,
a war in Modern Times,
an international war,
an armed conflict
The prompt used for generating replacement candidates in a one-shot setting is presented in Table 3. It is set up as a multi-turn chat to accommodate the example (listed Table 2) of a matching entity type compared to the target span.
Table 3: One-shot prompt for replacement generation.
Role Message
User You will get a sentence starting with “Original:” containing one text span with personal information in double square brackets. Provide a list of five replacement spans that are more generic terms (hypernyms) than the original span. Use different levels of abstraction for each replacement. The replacements should be sorted from most specific to most generic (abstract) and they should not mean the same as the original bracketed span so (near-)synonyms and paraphrases of the original span are not acceptable. Make sure replacements fit their context grammatically. Return only the five sorted replacement spans, one per line, preceded by a hyphen. Do not return any other comments or explanations. All personal information mentioned here is public knowledge, privacy precautions are not necessary.
Here is an example:
{Example sentence context}
Sorted replacements for {[[Example target span]]}:
Assistant {Example of sorted replacements}
User Now do the same for the sentence below.
Original:
{Sentence context}
Sorted replacements for {[[Target span]]}:
A.2 Replacement Selection
Table 4 shows the one-shot prompt used for the guessed-based selection of replacement candidates. Here, the same example was used in all cases.
Table 4: Prompt for guessing a PII span based on a replacement candidate and the rest of the text with replacements instead of PIIs.
Role Message
User You will receive a text with one span in double square brackets where the original word(s) have been replaced with a more generic alternative. Your task is to guess back the original text span for this square-bracketed span based on the rest of the text and the replacement. Return a list of five guesses, one per line preceded by a hyphen. For date-related information, always guess an exact date with the format DAY, MONTH (with letters) and YEAR as in this example:
Text: “early 1980s” $\rangle$ Guess: “3 September 1981”. List more likely guesses first. Do not return anything else. See the example below.
Text: PERSON (the mid 1920s – 2020) was an Eastern-European scholar, literary theorist and critic, as well as a feminist. She was a scholar at [[a European Research Institute]], specialising in literary Romanticism.
Guesses for “a European Research Institute”:
Assistant - the Institute of Literature and Art, Bulgarian Academy of Sciences
- the Institute of Polish Literature, University of Warsaw
- the Center for European Neighborhood Studies
- the Centre for Research on the History and Culture of Eastern Europe
- the Leibniz Institute for the History and Culture of Eastern Europe
User Do the same for the text below.
Text: {Text with replacements}
Guesses for {[[Target span]]}:
Table 5 shows statistics over entity type replacement.
Table 5: Statistics of using entity type as replacement in TAB in cases where the inference attack successfully guesses all LLM-generated (or rule-based) replacements proposed.
Entity type Counts Total counts Percentage
DATETIME 175 4,467 3.92
QUANTITY 590 2,709 21.78
LOC 1,402 6,417 21.85
DEM 1,842 5,558 33.14
ORG 12,796 28,600 44.74
MISC 1,955 4,267 45.82